# Coordinate In and Value Out: Training Flow Transformers in Ambient Space

### v1

@@ 4 Experiments
We evaluate ASFT on two challenging problems: image generation (FFHQ-256 (Karras et al., 2019), LSUN-Church-256 (Yu et al., 2015), ImageNet-128/256 (Russakovsky et al., 2015)) and 3D point cloud generation (ShapeNet (Chang et al., 2015) and Objaverse (Deitke et al., 2023)). Note that we use the same training recipe both tasks, adapted for changes in coordinate-value pair dimensions in different domains. See App. A for more implementation details and training settings.
ASFT enables practitioners to define the number of coordinate-value pairs to be decoded during training. In our experiments, we set the number of decoded coordinate-value pairs to 4096 for images with resolution 128$\times$128, 8192 for images with resolution 256$\times$256, and 2048 for point clouds unless mentioned otherwise. On image generation, we train models with small (S), base (B), large (L), and extra large (XL) sizes. For 3D point cloud generation we set the parameter count to match the model size in previous state-of-the-art approaches (i.e. LION (Vahdat et al., 2022)). Detailed configuration for all models can be found in Appendix A. During inference, we adopt black-box numerical ODE solver with maximal NFE as 100 for image generation (Song et al., 2021b) and an SDE sampler with 1000 steps for point cloud generation to match the settings in (Vahdat et al., 2022).
4.1 Image Generation in Function Space
Given that ASFT is a generative model for maps we compare it with other generative models of the same type, namely approaches that operate in function spaces. Tab. 1 shows a comparison of different image domain specific as well as function space models (e.g. approaches that model infinite-dimensional signals). ASFT surpasses other generative models in function space on both FFHQ (Karras et al., 2019) and LSUN-Church (Yu et al., 2015) at resolution $256\times 256$. Compared with generative models designed specifically for images, ASFT also achieves comparable or better performance. When scaling up the model size, ASFT-L demonstrates better performance than all the baselines on FFHQ-256 and Church-256, indicating that ASFT can benefit from increasing model sizes.
@@ 4.1 Image Generation in Function Space
ASFT-B (ours) 2.46 7.11
ASFT-L (ours) 2.18 5.51
@@
4.2 ImageNet
We also evaluate the performance of ASFT on large scale and challenging settings, we train ASFT on ImageNet at both 128$\times$128 and 256$\times$256 resolutions. On ImageNet-128, shown in Tab. 2, ASFT achieves an FID of $2.73$, which is a a competitive performance in comparison to diffusion or flow-based generative baselines including ADM (Dhariwal & Nichol, 2021), CDM (Ho et al., 2021), and RIN (Jabri et al., 2023) which use domain-specific architectures for image generation. Besides, comparing to PolyINR (Singh et al., 2023) which also operates on function space, ASFT achieves competitive FID, while obtaining better IS, precision and recall. The experimental results demonstrate the capabilities of ASFT in generating realistic samples on large scale datasets.
@@ 4.2 ImageNet
We report results of ASFT for ImageNet-256 on Tab. 3. Note that ASFT is slightly outperformed by latent space models like DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024). We highlight that these baselines rely on a pre-trained VAE compressor that was trained on datasets that are much larger than ImageNet, while ASFT was trained only with ImageNet data. In addition, ASFT achieves better performance than many of the baselines trained only with ImageNet data including ADM (Dhariwal & Nichol, 2021), CDM (Ho et al., 2021) and Simple Diffusion (U-Net) (Hoogeboom et al., 2023) which all use CNN-based architectures specific for image generation. Note that this is consistent with the results show in Tab. 1, where ASFT outperforms all function space approaches. When comparing with approaches using transformers architectures we find that ASFT obtains performance comparable to RIN (Jabri et al., 2022) and HDiT (Crowson et al., 2024), with slightly worse FID and slightly better IS. However, ASFT is a domain-agnostic architecture that can be trivially applied to different data domains like 3D point clouds (see Sect. 4.3 and 4.4). For completeness, we also include a comparison with very large U-Net transformer hybrid models, Simple Diffusion (U-ViT 2B) and VDM++ (U-ViT 2B) which both use approx. ×$\times 2.72$ more parameters than ASFT-XL, unsurprisingly, these much bigger capacity models outperform ASFT (see App. A for a more detailed comparison including training settings). We highlight that the simplicity of implementing and training ASFT models in practice, and the trivial extension to different data domains (as shown in Sect. 4.3) are strong arguments favouring our model. Finally, comparing with e.g. PolyINR (Singh et al., 2023) which is also a function space generative model we also find comparable performance, with slight worse FID but better Precision and Recall. It is worth noting that (Singh et al., 2023) applies a pre-trained DeiT model as the discriminator (Singh et al., 2023). Whereas our ASFT makes no such assumption about the function or pre-trained models, enabling to trivially apply ASFT to other domains like 3D point clouds (see Sect. 4.3).
To demonstrate the scalability of ASFT we train models of different sizes including small (S), base (B), large (L), and extra-large (XL) on ImageNet-256. We show the performance of different model sizes using FID-50K in Fig. 3(a). We observe a clear improving trend when increasing the number of parameters as well as increasing training steps. This demonstrates that scaling the total training Gflops is important to improved generative results as in other ViT-based generative models (Peebles & Xie, 2023; Ma et al., 2024). Due to the flexibility of cross-attention decoder in ASFT, one can easily conduct random sub-sampling to reduce the number of decoded coordinate-value pairs during training which significantly saves computation. Fig. 3(b) shows how number of decoded coordinate-value pairs affects the model performance as well as Gflops in training. An image of resolution 256$\times$256 contains 65536 pixels in total which is the maximal number of coordinate-value pairs during training. As see in Fig. 3(b), a model decoding 4096 coordinate-value pairs saves more than 20% Gflops over one decoding 16384. This provides us with an effective training recipe, which saves computation by only decoding a subset of $12\%$ of the image pixels during training. Interestingly, we see a performance drop when densely decoding 16384 coordinate value pairs. We hypothesize this could be due to optimization challenges of decoding large numbers of pairs and leave further analysis for future work.
@@
BigGAN-deep (Brock et al., 2019) 6.95 171.4 0.87 0.28
PolyINR (Singh et al., 2023) 2.86 241.4 0.71 0.39
@@
ADM (Dhariwal & Nichol, 2021) 10.94 100.9 0.69 0.63
CDM (Ho et al., 2021) 4.88 158.7 - -
Simple Diff. (U-Net) (Hoogeboom et al., 2023) 3.76 171.6 - -
RIN (Jabri et al., 2023) 3.42 182.0 - -
HDiT (cfg=1.3) (Crowson et al., 2024) 3.21 220.6 - -
Simple Diff. (U-ViT) Hoogeboom et al. (2023) 2.77 211.8 - -
VDM++ (U-ViT)  (Kingma & Gao, 2023) 2.12 267.7 - -
ASFT-XL (ours) (cfg=1.5) 3.74 228.8 0.82 0.52
@@
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) FID-50K over training iterations with different model sizes, where we see clear benefits of scaling up model sizes. (b) FID-50K over training iterations with different number of decoded coordinate-value pairs during training and the corresponding compute cost for a single forward pass.
4.3 ShapeNet
To show the domain-agnostic prowess of ASFT we also tackle 3D point cloud generation on ShapeNet (Chang et al., 2015). Note that our model does not require training separate VAEs for point clouds, tuning their corresponding hyper-parameters or designing domain specific networks. We simply adapt our architecture for the change in dimensionality of coordinate-value pairs (e.g. $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ for images to $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ for 3D point clouds.). Note that for 3D point clouds, the coordinates and values are equivalent. In this setting, we compare baselines including LION (Vahdat et al., 2022) which is a recent state-of-the-art approach that models 3D point clouds using a latent diffusion type of approach. Following Vahdat et al. (2022) we report MMD, COV and 1-NNA as metrics. To have a straightforward comparison with baselines, we train ASFT-B with to approximately match the number of parameters as LION (Vahdat et al., 2022) (110M for LION vs 108M for ASFT) on the same datasets (using per sample normalization as in Tab. 17 in Vahdat et al. (2022)). We show results for category specific models and for an unconditional model jointly trained on 55 ShapeNet categories in Tab. 4. ASFT-B obtains strong generation results on ShapeNet despite being a domain agnostic approach and outperforms LION in most datasets and metrics. Note that ASFT-B has comparable number of parameters and the same inference settings than LION so this is fair comparison. Finally, we also report results for a larger model ASFT-L (with ×$\times 2$ the parameter count as LION) to investigate how ASFT improves as with increasing model size. We observe that with increasing model size, ASFT typically achieves better performance than the base version. This further demonstrates scalability of our model on ambient space of different data domains.
[…]
Category Model CD EMD CD EMD CD EMD
ShapeGF (Cai et al., 2020) 0.3130 0.6365 45.19 40.25 81.23 80.86
SP-GAN (Li et al., 2021) 0.4035 0.7658 26.42 24.44 94.69 93.95
ASFT-B (ours) (108M) 0.2861 0.5156 43.38 47.54 75.55 64.95
ASFT-L (ours) 0.2880 0.5052 44.44 47.16 62.20 62.96
ShapeGF (Cai et al., 2020) 3.7243 2.3944 48.34 44.26 58.01 61.25
SP-GAN (Li et al., 2021) 4.2084 2.6202 40.03 32.93 72.58 83.69
Chair GCA (Zhang et al., 2021) 4.4035 2.5820 45.92 47.89 64.27 64.50
ASFT-B (ours) (108M) 3.6310 2.1725 46.67 53.31 55.43 51.13
ASFT-L (ours) 3.5145 2.1860 49.39 49.84 50.52 51.66
ShapeGF (Cai et al., 2020) 1.0200 0.8239 44.03 47.16 61.79 57.24
SP-GAN (Li et al., 2021) 1.1676 1.0211 34.94 31.82 87.36 85.94
Car GCA (Zhang et al., 2021) 1.0744 0.8666 42.05 48.58 70.45 64.20
ASFT-B (ours) (108M) 0.9923 0.7692 43.46 47.44 60.36 53.27
ASFT-L (ours) 0.9660 0.7846 44.03 48.86 53.83 54.55
All (55 cat) ASFT-B (ours) (108M) 3.2586 2.1328 49.00 50.40 54.65 55.70
ASFT-L (ours) 3.1775 1.9794 49.80 52.39 51.80 53.90
4.4 Objaverse
We also experiment ASFT on large-scale 3D point generation conditioned on images. To this end, we train an image-to-point-cloud generation model on Objaverse (Deitke et al., 2023), which contains 800k 3D objects of great variety. In particular, conditional information (i.e., an image) is integrated to our model through cross-attention. For each object in Objaverse, we sample point cloud with $16k$ points. To get images for conditioning, each object is rendered with 40 degrees field of view, $448\times 448$ resolution, at 3.5 units on the opposite sides of $x$ and $z$ axes looking at the origin. We extract features via DINOv2 (Oquab et al., 2023) which is concatenated with Plucker ray embedding (Plucker, 2018) of each patch in DINOv2 feature. In each block, the learnable latent vector $z_{f_{t}}$ cross attends to image feature.
@@ 4.4 Objaverse
Tab. 5 lists the performance of ASFT in comparison of recent baselines on Objaverse. We report ULIP-I (Xue et al., 2024) and P-FID (Nichol et al., 2022) following CLAY (Zhang et al., 2024). PointNet++ (Qi et al., 2017a; b; Nichol et al., 2022) is employed to evaluate P-FID. ULIP-I is an analogy to CLIP for text-to-image generation. ULIP-I is measured as the cosine similarity between point-cloud features from ULIP-2 model (Xue et al., 2024) and image features from CLIP model (Radford et al., 2021). Numbers of baseline models are directly borrowed from CLAY (Zhang et al., 2024). We calculate the metrics of our ASFT on 10k sampled point clouds. In our case, P-FID is measured on point clouds with 4096 points following Shape-E (Jun & Nichol, 2023) while ULIP-I is measured on point clouds with 10k points following ULIP-2 (Xue et al., 2024). Note that since CLAY (Zhang et al., 2024) is not open-source, we do not have the access to the exact evaluation setting or the conditional images rendered from Objaverse. But all evaluation settings of ASFT are provided for reproduction purpose. As shown in Tab. 5, our ASFT achieves strong performance on large-scaled image-conditioned 3D generative tasks. Compared with CLAY (Zhang et al., 2024), which is a 2-stage latent diffusion model, ASFT demonstrates very strong performance on both ULIP-I and P-FID. Fig. 9 and 10 show additional examples of sampled point clouds and corresponding conditional images. As shown, ASFT learns to generate 3D objects with rich details that match the conditional images ultimately being able to generate a continuous surface.
@@ 4.5 Resolution Agnostic Generation
An interesting property of ASFT is that it decodes each coordinate-value pair independently, allowing resolution to change during inference. At inference the user can define as many coordinate-value pairs as desired where the initial value of each pair at $t=1$ is drawn from a Gaussian distribution. We show qualitative results of resolution agnostic generation for both images and point clouds. Fig. 4(a) show images sampled at resolution $512^{2}$, $1024^{2}$, and $2048^{2}$ (together with their $256^{2}$ resolution counterparts generated from the same seed) from ASFT trained on ImageNet-256. Even though the model has not been trained with any samples at higher resolutions, it can still generate realistic images with high-frequency details. Fig. 4(b) shows point cloud with $32k$ and $128k$ points from ASFT trained on Objaverse with only $16k$ points points per sample (we visualize the generated $16k$ point could generated from the same seed). Similarly, ASFT generates dense and realistic point cloud in 3D without actually being trained on such high density points. These results show that ASFT is not trivially overfitting to the training set of points but rather learning a continuous density field in 3D space from which an infinite number of points could be sampled. Generally speaking, this also provides the potential to efficiently train flow matching generative models without the need to use large amounts of expensive high resolution data, which can be hard to collect in data domains other than images.
Figure 4: (a) Images generated at $256^{2}$, $512^{2}$, $1024^{2}$, and $2048^{2}$ resolution from an ASFT trained on ImageNet-256. (b) 3D point clouds generated by ASFT containing $16k$, $32k$, and $128k$ points from an ASFT trained on Objaverse with $16k$ points (conditioned image shown in bottom left).

### v2

@@ 4 Experiments
We evaluate INRFlow on two challenging problems: image generation (FFHQ-256 (Karras et al., 2019), LSUN-Church-256 (Yu et al., 2015), ImageNet-128/256 (Russakovsky et al., 2015)), image-to-3D point cloud generation (Objaverse (Deitke et al., 2023)) and protein folding (SwissProt (Boeckmann et al., 2003)). Note that we use the same training recipe for all tasks, adapted for changes in coordinate-value pair dimensions in different domains. See App. A for more implementation details and training settings.
4.1 Image Generation
Given that INRFlow is an probabilistic extension of INRs we compare it with other generative models of the same type, namely approaches that operate in continuous function spaces. Tab. 1 shows a comparison of different image domain specific, as well as, function space models (e.g. approaches that model infinite-dimensional signals). INRFlow surpasses other generative models in function space on both FFHQ (Karras et al., 2019) and LSUN-Church (Yu et al., 2015) at resolution $256\times 256$. Compared with generative models designed specifically for images, INRFlow also achieves comparable or better performance. When scaling up the model size, INRFlow-L demonstrates better performance than all the baselines on FFHQ-256 and Church-256, indicating that INRFlow can benefit from increasing model sizes.
@@ 4.1 Image Generation
INRFlow-B (ours) 2.46 7.11
INRFlow-L (ours) 2.18 5.51
@@
We also evaluate the performance of INRFlow on large scale settings previously untapped for domain agnostic approaches, training INRFlow on ImageNet at both 128$\times$128 and 256$\times$256 resolutions. On ImageNet-128, shown in Tab. 2, INRFlow achieves an FID of $2.73$, which is a a competitive performance in comparison to diffusion or flow-based generative baselines including ADM (Dhariwal & Nichol, 2021), CDM (Ho et al., 2021), and RIN (Jabri et al., 2023) which use domain-specific architectures for image generation. Besides, comparing to PolyINR (Singh et al., 2023) which also operates on function space, INRFlow achieves competitive FID, while obtaining better IS, precision and recall. In addition, we report results of INRFlow for ImageNet-256 on Tab. 3. We observed that INRFlow is slightly outperformed by latent space models like DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024). We highlight that these baselines rely on a pre-trained VAE compressor that was trained on datasets (i.e. 9.29M images) that are much larger than ImageNet, while INRFlow was trained only with ImageNet data. In addition, INRFlow achieves better performance than many of the baselines trained only with ImageNet data including ADM (Dhariwal & Nichol, 2021), CDM (Ho et al., 2021) and Simple Diffusion (U-Net) (Hoogeboom et al., 2023) which all use CNN-based architectures specific for image generation. Note that this is consistent with the results show in Tab. 1, where INRFlow outperforms all function space approaches.
When comparing with approaches using transformer architectures we find that INRFlow obtains performance comparable to RIN (Jabri et al., 2022) and HDiT (Crowson et al., 2024), with slightly worse FID and slightly better IS. However, INRFlow is a domain-agnostic architecture that can be trivially applied to different data domains like 3D point clouds or protein structure data (see Sect. 4.2 and Sect. 4.3). For completeness, we also include billion scale U-Net transformer hybrid models, Simple Diffusion (U-ViT 2B) and VDM++ (U-ViT 2B). We highlight that the simplicity of implementing and training INRFlow models in practice, and the trivial extension to different data domains are strong arguments favoring INRFlow. Finally, comparing with e.g. PolyINR (Singh et al., 2023) which is also a function space generative model we also find comparable performance, with slight worse FID but better Precision and Recall. It is worth noting that (Singh et al., 2023) applies a pre-trained DeiT model as the discriminator (Singh et al., 2023). Whereas our INRFlow makes no such assumption about the function or pre-trained models, enabling to trivially apply INRFlow to other domains like 3D point clouds (see Sect. 4.2) or protein folding.
@@
BigGAN-deep (Brock et al., 2019) 1.28M - - 6.95 171.4 0.87 0.28
PolyINR (Singh et al., 2023) 1.28M - - 2.86 241.4 0.71 0.39
@@
ADM (Dhariwal & Nichol, 2021) 1.28M 554M 507M 10.94 100.9 0.69 0.63
CDM (Ho et al., 2021) 1.28M - - 4.88 158.7 - -
Simple Diff. (U-Net) (Hoogeboom et al., 2023) 1.28M - - 3.76 171.6 - -
RIN (Jabri et al., 2023) 1.28M 410M 614M 3.42 182.0 - -
HDiT (cfg=1.3) (Crowson et al., 2024) 1.28M 557M 742M 3.21 220.6 - -
Simple Diff. (U-ViT) (Hoogeboom et al., 2023) 1.28M 2B 1.4B 2.77 211.8 - -
VDM++ (U-ViT)  (Kingma & Gao, 2023) 1.28M 2B 1.4B 2.12 267.7 - -
INRFlow-XL (ours) (cfg=1.5) 1.28M 733M 780M 3.74 228.8 0.82 0.52
@@
4.2 Image-to-3D
We also showcase that INRFlow can directly integrate conditional information like images. We train an image-to-point-cloud INRFlow model on Objaverse (Deitke et al., 2023), which contains 800k 3D objects of wide variety, to illustrate the capability of INRFlow on large-scale 3D generative tasks. In particular, conditional information (i.e., an image) is integrated to our model through cross-attention. We train INRFlow with batch size 384 for 500k iterations. During sampling, we use an Euler-Maruyama sampler (Ma et al., 2024) with 500 steps to generate point clouds (see App. A.2 for more details on architecture and objaverse data generation).
@@ 4.2 Image-to-3D
Tab. 4 shows the performance of INRFlow in comparison with recent baselines on Objaverse. We report ULIP-I (Xue et al., 2024) and P-FID (Nichol et al., 2022) following CLAY (Zhang et al., 2024). PointNet++ (Qi et al., 2017a, b; Nichol et al., 2022) is employed to evaluate P-FID. ULIP-I is an analogy to CLIP for text-to-image generation. ULIP-I is measured as the cosine similarity between point-cloud features from ULIP-2 model (Xue et al., 2024) and image features from CLIP model (Radford et al., 2021). The performance numbers of baseline models are directly borrowed from CLAY (Zhang et al., 2024). We calculate the metrics of our INRFlow on 10k sampled point clouds. In our case, P-FID is measured on point clouds with 4096 points following Shape-E (Jun & Nichol, 2023) while ULIP-I is measured on point clouds with 10k points following ULIP-2 (Xue et al., 2024). Note that since CLAY (Zhang et al., 2024) is not open-source, we do not have the access to the exact evaluation setting or the conditional images rendered from Objaverse. But all evaluation settings of INRFlow are provided for reproduction purpose. As shown in Tab. 4, our INRFlow achieves strong performance on image-to-3D generative tasks. Compared to CLAY (Zhang et al., 2024), which is a 2-stage latent diffusion model, INRFlow demonstrates very strong performance on both ULIP-I and P-FID.
@@
Figure 3: Examples of protein structures predicted by INRFlow on SwissProt, together with their LDDT and TM scores. The GT structures are depicted in green while the generated structures are show in orange. INRFlow accurately captures the global spatial distribution of protein backbones generating reasonable 3D structures for different protein sequences.
Fig. 4(b) show examples of sampled point clouds and corresponding conditional images. As discussed in §4.4, INRFlow trained on Objaverse also enjoys the flexibility of resolution agnostic generation. In the App, Fig. 12 we show additional results sampled with more points than what the model was trained on. As shown, INRFlow learns to generate 3D objects with rich details that match the conditional images ultimately being able to generate a continuous surface.
4.3 Protein Folding
We now showcase the domain-agnostic prowess of INRFlow applying it to the protein folding problem (Jumper et al., 2021). From an ML perspective, this problem is a conditional 3D generation problem where we are given the amino-acid sequence (e.g. a sequence of discrete symbols from a vocabulary of 20 possible amino-acids) and we need to generate a 3D coordinate for each atom in the protein, where different amino-acids can have different numbers of atoms. In our experiments we use SwissProt set (Boeckmann et al., 2003) taking the ground truth structures from the AlphaFold Database (Varadi et al., 2022).
We adapt the coordinate and signal in INRFlow so that the coordinate becomes a feature representation of each atom in the protein and the signal represents the 3D point for that particular atom. This is similar to recent work on protein folding and conformer generation, which directly predicts atom coordinates as opposed to frames of reference (Wang et al., ; Abramson et al., 2024). For the spatially-aware latents we aggregate information from all the atoms corresponding to the same amino-acid into a single latent vector (see App. A.3 for more details on architecture and training recipe).
We show qualitative results on Fig. 3 and additional results on Fig. 13. INRFlow accurately captures the global distribution of protein backbones generating reasonable 3D structures for different protein sequences. These initial results are very encouraging and open up exciting design spaces for future work developing protein folding network architectures that can leverage general purpose transformer blocks.
@@ 4.4 Resolution Agnostic Generation
Figure 4: Examples of resolution agnostic generation for INRFlow models trained on ImageNet-256 (a), and Objaverse-16k in (b). To generate samples at higher resolutions than the one in training we fix the initial noise seed and increase the number of coordinate-value pairs evaluated by the model. Even though INRFlow was only train with samples at a fixed resolution (256 for ImageNet and 16k for Objaverse), it can still generate realistic samples at higher resolutions. These results show that INRFlow is learning a continuous probability density field.
An interesting property of INRFlow is that it decodes each coordinate-value pair independently, allowing resolution changes of resolution for generation. At inference time the user can define as many coordinate-value pairs as desired where the initial value of each pair at $t=1$ is drawn from a Gaussian distribution. We now quantitatively evaluate the performance of INRFlow in this setting. In Tab. 5 we compare the FID of different recipes. First, INRFlow is trained on FFHQ-256 and bilinear or bicubic interpolation is applied to the generated samples to get images at 512. On the other hand, INRFlow can directly generate images at resolution 512 by simply increasing the number of coordinate-value pairs during inference without further tuning. As shown in Tab. 5 , INRFlow achieves lower FID when compared with other manually designed interpolation methods, showcasing the benefit of developing generative models on ambient space.
INRFlow Bilinear Bicubic
FID($\downarrow$) 23.09 35.05 24.34
Table 5: FID of different super resolution methods to generate images at resolution $512\times 512$ for INRFlow trained on FFHQ-256.
We show examples of resolution agnostic generation for INRFlow models trained on ImageNet-256 Fig. 4(a) and Objaverse-16k in Fig. 4(b). Even though INRFlow was only train with samples at a fixed resolution (256 for ImageNet and 16k for Objaverse), it can still generate realistic samples at higher resolutions. For example, Fig.  4(a) shows samples generated from INRFlow trained on ImageNet-256 and sampled at resolutions up to 2k (see additional examples in Fig. 14). Fig. 4(b) shows point cloud with up to 128k points from INRFlow trained on Objaverse with only 16k points points per sample. It is worth noting that in INRFlow there’s no cascading (Ho et al., 2022b) or multiple resolution training (Gu et al., 2023), we simply fix the initial noise seed and increase number of coordinate-value pairs that are evaluated. These results show that INRFlow is not trivially overfitting to the coordinate-value pairs in the training but rather learning a continuous probability density field in space from which an infinite number of points could be sampled. Generally speaking, this also provides the potential to efficiently train flow matching generative models without the need to use large amounts of expensive high resolution data, which can be hard to collect in some domains.